vibesss2x:

At a farmer’s market, a basket of strawberries costs $20 and a basket of peaches costs $18. On Saturday, $430 worth of strawberries and peaches were sold, and a total of 23 baskets were sold. If x represents baskets of strawberries and y represents baskets of peaches, how many baskets of peaches were sold on Saturday?

Mercury:

"x = baskets of strawberries" "y = baskets of peaches" total baskets = 23 so x + y = 23 "430 worth of produce" the cost of produce can be found by taking (cost of strawberries)(# of baskets of strawberries) + (cost of peaches)(# of baskets of peaches) = total cost therefore 20x + 18y = 430

Mercury:

solve the system: 20x + 18y = 430 x + y = 23 by elimination, substitution, or graphing